OpenStudy (anonymous):

Help me pleaseee !!! How can the properties of exponents help solve exponential equations?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

just showing some rules.... \(\Large\color{blue}{ a^{b} \div a^{c}=a^{b-c} }\) \(\Large\color{blue}{ a^{b} \times a^{c}=a^{b+c} }\) \(\Large\color{blue}{ (a^{b})^c=a^{b \times c} }\) And mainly, \(\Large\color{red}{ \bf a^{-b}=\frac{\LARGE{1}}{\LARGE{a^b}} }\) (where ` b ` is a positive number)

OpenStudy (solomonzelman):

These Properties, can be used to simplify the exponential equations, very easily.

OpenStudy (anonymous):

Thank you so much i couldnt even begin to express my gratitude these questions are worded badly and hard to understand